Circular No. 7083

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999A IN NGC 5874 M. Modjaz, E. Halderson, T. Shefler, J. Y. King, M. Papenkova, W.

Li, R. R. Treffers, and A. V. Filippenko, University of California at Berkeley, report their discovery of an apparent supernova during the course of the Lick Observatory Supernova Search (cf. IAUC 6627) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 1999A was discovered and confirmed on unfiltered observations taken on Jan. 10.6 (mag ∼ 18.3) and 11.6 UT (mag ∼ 18.2; limiting mag 19.5–20.0). The new object is located at α = 15^{h}07^{m}55^{s}.28, δ = +54^{o}45′40′′.0 (equinox 2000.0), which is ∼ 29′′.9 east and 30′′.1 north of the nucleus of NGC 5874. There is a star of mag 16.2 very close to SN 1999A with position end figures 55^{s}.10, 45′′.2. A KAIT image of the same field on 1998 July 23.3 (limiting mag ∼ 18.5) shows nothing at the position of SN 1999A, nor do the first and second Digital Sky Survey images of the field (limiting mag > 20).

SUPERNOVA 1998co IN NGC 7131 E. Prosperi, Larciano, Italy, provides the following position for SN

1998co from a Mount Hopkins 1.2-m telescope image obtained on 1998 June

22 (see http://cfa-www.harvard.edu/cfa/oir/Research/supernova/

RecentSN.html): α = 21^{h}47^{m}36^{s}.45, δ = −13^{o}10′42′′.3 (equinox 2000.0).

COMET C/1995 O1 (HALE-BOPP) Ephemeris extension to IAUC 7047 (orbital elements MPC 32410):

1999 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Jan. 2 5^{h}48.^{m}1 −73^{o}47.^{′}0 7.394 7.335 82.^{o}7 7.^{o}6 11.0
12 5 29.7 −73 26.9 7.479 7.418 82.7 7.6 11.1
22 5 14.2 −72 50.1 7.563 7.501 82.6 7.5 11.1
Feb. 1 5 02.1 −72 00.6 7.647 7.583 82.6 7.4 11.2
11 4 53.7 −71 02.5 7.729 7.664 82.6 7.3 11.3
21 4 48.59 −69 59.6 7.810 7.746 82.7 7.3 11.4
Mar. 3 4 46.46 −68 55.1 7.889 7.827 82.8 7.2 11.4
13 4 46.83 −67 51.6 7.967 7.907 83.0 7.2 11.5
23 4 49.28 −66 51.2 8.044 7.987 83.2 7.1 11.6
Apr. 2 4 53.44 −65 55.6 8.119 8.067 83.5 7.1 11.6

Visual m_{1} estimates: Jan. 1.79 UT, 11.2 (A. Pearce, Nedlands, W. Australia, 0.41-m reflector); 5.41, 10.2 (S. T. Rae, Te Awamutu, New Zealand, 0.11-m reflector); 7.42, 10.4 (Rae); 10.65, 11.0 (Pearce).

1999 January 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green